a sphere whose radius at time t is b+acosnt is surrounded by liquid extending to infinity under no force . prove that the pressure at distance r from the centre is less than the pressure at an infinite distance by pn^2/r(b+acosnt)[a(1-3sin^2 nt)+bcos nt+a/2r^3 (sin^2 nt)(b+acos nt)^3]


1
Expert's answer
2021-06-24T11:59:40-0400

Given,

The radius of the sphere at the time t is "(r)=b+a\cos(nt)"

"dr=-an\sin(nt)"

Let the uniform density of the sphere be "\rho"

"dV = A.dr"

So, mass of the partial sphere "(dm)=\rho A.dr"

Weight of the sphere "(W)=\rho g A. dr"

"g=\frac{GM}{r^2}=\frac{4\pi Gr^3}{3r^2}=\frac{4}{3}\pi Gr"

Let the pressure at the radius r is "P"

So, pressure at "r+dr" be "P+dP"

Now, equating the downward force and upward force

"A(P+dP)+\frac{4\pi A G\rho^2 rdr}{3}-\frac{4\pi r^3}{3}\rho\times\frac{4}{3}\pi Gr=AP"


"\Rightarrow (dP)+\frac{4\pi G\rho^2 rdr}{3}-\frac{4\pi r^3}{3A}\rho\times\frac{4}{3}\pi Gr=0"


"\Rightarrow" "dP=\frac{4\pi r^3}{3A}\rho\times\frac{4}{3}\pi Gr-\frac{4\pi G\rho^2 rdr)}{3}"


"\Rightarrow dP=\frac{4\pi (b+a\cos(nt))^4}{3A}\rho\times\frac{4}{3}\pi G-\frac{4\pi G\rho^2 (b+a\cos(nt))(-an\sin(nt)}{3}"